OpenStudy (anonymous):

can't find the integral of this function v(t) = (1+t^2)^(1/3) anyone know how?

OpenStudy (anonymous):

\[\int\limits \sqrt[3]{1 + t^2}\] this is what i cant find

OpenStudy (idku):

\[\large \int\limits_{ }^{ }\sqrt[3]{1+t^2}dt\]

OpenStudy (anonymous):

yep thats right

OpenStudy (idku):

well, seems as though it doesn't have closed form based on my first impression. Lets set u=1+t² and see wat we get, did you try that?

OpenStudy (anonymous):

yep, didn't get anywhere. be my guest though im pretty bad at u substitutions :P

OpenStudy (idku):

oh, my trig sub...

OpenStudy (idku):

\(\large \displaystyle\int\limits_{ }^{ }\sqrt[3]{1+t^2}dt\) \(\large \displaystyle t=\tan\theta \)

OpenStudy (anonymous):

trig sub...? never heard of it.

OpenStudy (idku):

lets see what then, u=tan(theta) du=sec^2theta \(\large \displaystyle\int\limits_{ }^{ }\sec^2\theta\sqrt[3]{1+\tan^2\theta}~d\theta \) \(\large \displaystyle\int\limits_{ }^{ }\sec^{2+2/3}\theta\) that is bullsh... as well

OpenStudy (anonymous):

wolfram is giving me something i dont even understand .

OpenStudy (idku):

no simple function answer

OpenStudy (idku):

are you sure you aren't given limits of integration, and this is not a riemman sum type of question?

OpenStudy (idku):

or maybe it is sort of \(\large \displaystyle f(x)=\int\limits_{2 }^{ \cos(x)}\sqrt[3]{1+t^2}dt\) and you need f'(x) ?

OpenStudy (anonymous):

this is my given question

OpenStudy (idku):

this is much better

OpenStudy (idku):

can I use y instead of x (as y(t) ?)

OpenStudy (idku):

You can use a stepsize on that one I think

OpenStudy (anonymous):

interesting i figured i just had to find s(t) and just change the constant and plug in t = 3

OpenStudy (idku):

well, that is the task as far as the steps go, but s(t)=?

OpenStudy (idku):

that is the question where we don't arrive at a simple conclusion just like this, right?

OpenStudy (idku):

have you heard of a step-size, Euler's method or anything like this?

OpenStudy (anonymous):

no sorry :(

OpenStudy (idku):

\(\large \displaystyle f'(t)=\sqrt[3]{1+t^2}\) y(0)=2, find y(3)=? lets use a step-size of h=1 for now (but with h=1/2 or even less you can get a better approximation. Each step of size h is as follows: \(\large \displaystyle \left(x_n,y_n\right)=\left(x_{n-1}+h,~y_{n-1}~+h\cdot \color{red}{f'(x_{n-1},y_{n-1})}~\right)\)

OpenStudy (idku):

maybe we won't understand why it is like this right now, but you can see the formula, without completely going like wt(f)....

OpenStudy (idku):

\(\large \displaystyle \left(t_n,y_n\right)=\left(t_{n-1}+h,~y_{n-1}~+h\cdot \color{red}{f'(t_{n-1},y_{n-1})}~\right)\) excuse me it should be t.

OpenStudy (idku):

\(\large \displaystyle \left(x_n,y_n\right)=\left(x_{n-1}+h,~y_{n-1}~+h\cdot \color{red}{f'(x_{n-1},y_{n-1})}~\right)\) you are given: \(\large \displaystyle f'(t)=\sqrt[3]{1+t^2}\) and a first point of (t=0, y=2)

OpenStudy (idku):

\(\large \displaystyle \left(x_1,y_1\right)=\left(x_{0}+1,~2~+1\cdot \color{red}{f'(0,2)}~\right)\) this would be the first step starting from (0,2), and h=1

OpenStudy (idku):

are you getting it at least a little? (sorry I am bad at explaining)

OpenStudy (idku):

ok, I will give you the entire process wth h=1, and then you will tell me how much you get out of it.

OpenStudy (anonymous):

lol sort of , i'm not quite sure what the h represents

OpenStudy (idku):

you are starting from (0,2) right? that is the first given point  so it is the size of the step, in other words, we are carefully drawing the approximation for the graph of y(t) based on y'(t) taking it with little stpes of size h (of size 1 in this case)

OpenStudy (idku):

we don't know where this step size of 1 would land, and usiong this formula we can find that out.

OpenStudy (idku):

and from there, when we do the step and find the next point, we will take another step... ultimately, we wil get to the y(3)

OpenStudy (idku):

\(\large \displaystyle \left(x_n,y_n\right)=\left(x_{n-1}+h,~y_{n-1}~+h\cdot \color{red}{f'(x_{n-1},y_{n-1})}~\right)\) \(\large \displaystyle \left(x_1,y_1\right)=\left(x_{0}+1,~2~+1\cdot \color{red}{f'(0,2)}~\right)\) (there just isn't a y in this case, so would just plug in 0 for t) I haven't ever done Euler's for explicitly defined functions. \(\large \displaystyle \left(x_1,y_1\right)=\left(x_{0}+1,~2~+1\cdot \color{red}{\sqrt[3]{1+0^2}}~\right)\)

OpenStudy (idku):

oh, you know x_0 is 0 , i forgot to put that in

OpenStudy (anonymous):

okay wait one sec , why are we even finding (x1,y1)?

OpenStudy (idku):

\(\large \displaystyle \left(x_1,y_1\right)=\left(0+1,~2~+1\cdot \color{red}{\sqrt[3]{1+0^2}}~\right)\) \(\large \displaystyle \left(x_1,y_1\right)=\left(1,~3\right)\)

OpenStudy (idku):

because that is the next point

OpenStudy (idku):